OpenStudy (anonymous):

Find each square 1) (w-12)^2 2)(b-5)^2 3)(6x-8)^2 4)(9j-2)^2 PLEASE HELP :)

OpenStudy (mathstudent55):

All of them follow the same pattern. (a - b)^2 = a^2 - 2ab + b^2 I'll do the first one as an example. In your first problem above you have (w - 12)^2. You are squaring w - 12. The first term of w - 12 is w, and the second term is -12. Square the first term and place it on the left. w^2 Then take the first term times the second term times 2 and write it down: w * (-12) * 2 = -24w, so now you have w^2 - 24w Then square the second term and write it down: w^2 - 24w + 144 That is it. Now try the others. If you need help, ask, but try it first.

OpenStudy (anonymous):

Thank you the expalining was great -)